OpenStudy (anonymous):

Solve the following system of equations. Please show your work to receive full credit. x - y = 13 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

hint: add bouth equations

OpenStudy (dan815):

basically you are trying to get 1 equation with just 1 unknown so u can solve it

OpenStudy (dan815):

the first method you always learn is just substitution, so say you arrange hte first equation x-y =13 x-13=y no sub this y into the 2nd equation and you will have just 1 uknown 2x+y=2 2x+(x-13)=2 simplify and solve for x 2x+x-13=2 3x=15 x=5

OpenStudy (dan815):

now from your initial relationship x-13=y x=5 so, 5-13=y =?

OpenStudy (anonymous):

The easiest way out in this specific problem is by adding the two equation... When you add the two equation...'y' eliminates to give the equation 3x=15 and therefor x=5! Since the first equation is less complicated when compared to second....substitute x=3 in the first equation and get the value of y.... That is y=3-13=-10